# From Ambiguity to Verdict: A Semiotic-Grounded Multi-Perspective Agent for LLM Logical Reasoning

### v1

@@ Abstract
Logical reasoning is a fundamental capability of large language models (LLMs). However, existing studies largely overlook the interplay between logical complexity and semantic complexity, resulting in methods that struggle to address challenging scenarios involving abstract propositions, ambiguous contexts, and conflicting stances, which are central to human reasoning. For this gap, we propose LogicAgent, a semiotic-square–guided framework designed to jointly address logical complexity and semantic complexity. LogicAgent explicitly performs multi-perspective deduction in first-order logic (FOL), while mitigating vacuous reasoning through existential import checks that incorporate a three-valued decision scheme (True, False, Uncertain) to handle boundary cases more faithfully. Furthermore, to overcome the semantic simplicity and low logical complexity of existing datasets, we introduce RepublicQA, a benchmark that reaches college-level difficulty (FKGL = 11.94) and exhibits substantially greater lexical and structural diversity than prior benchmarks. RepublicQA is grounded in philosophical concepts, featuring abstract propositions and systematically organized contrary and contradictory relations, making it the most semantically rich resource for evaluating logical reasoning. Experiments demonstrate that LogicAgent achieves state-of-the-art performance on RepublicQA, with a 6.25% average gain over strong baselines, and generalizes effectively to mainstream logical reasoning benchmarks including ProntoQA, ProofWriter, FOLIO, and ProverQA, achieving an additional 7.05% average gain. These results highlight the strong effectiveness of our semiotic-grounded multi-perspective reasoning in boosting LLMs’ logical performance.
@@ 1 Introduction
Logical reasoning Smith (2003) plays a central role in human cognition, enabling structured transitions from ambiguous inputs to definitive conclusions. In AI Cohen et al. (2020); Vaswani et al. (2017), it underpins tasks such as commonsense reasoning Wang et al. (2024), mathematical proof Wang et al. (2023); Eisner et al. (2024); Gao et al. (2023), and philosophical thinking Paul (2013). However, robust logical reasoning in natural language remains challenging due to (1) semantic complexity Tuggy (1993), where expressions admit multiple interpretations or surface forms, and (2) logical complexity Gibson (1998), which requires reasoning over contextual premises and semantic interactions Zhang et al. (2020); Ding et al. (2024).
Recent approaches such as Chain-of-Thought (CoT) prompting Wei et al. (2022), Tree-of-Thought Yao et al. (2023), and Least-to-Most Zhou et al. (2022) have improved step-by-step reasoning, while symbolic-enhanced frameworks Pan et al. (2023); Ryu et al. (2024); Yang et al. (2023); Qi et al. (2025) further combine large language models (LLMs) Achiam et al. (2023); Guo et al. (2025); Zhao et al. (2023b) with first-order logic (FOL) solvers. These methods achieve strong results on widely used benchmarks such as ProntoQA Saparov & He (2022), ProofWriter Tafjord et al. (2020), FOLIO Han et al. (2022) and ProverQA Qi et al. (2025). Nevertheless, existing benchmarks remain confined to relatively simple and determinate semantics, often centered on everyday scenarios with shallow relations or logic problems that lack intricate semantic structure. Consequently, current evaluations fail to capture the full complexity of real-world reasoning, where propositions frequently involve philosophical abstraction, opposing stances, and deeper latent logical relations.
@@
To capture the complex semantics and deep logical relations, we draw inspiration from Greimas’ Semiotic Square Greimas et al. (1982); Greimas (1987; 1988), a structuralist framework that extends binary oppositions into a four-part structure. It encompasses both contraries (e.g., $S_{1}$ vs. $S_{2}$, which cannot both be true but may both be false under non-empty domains) and contradictions (e.g., $S_{1}$ vs. $\lnot S_{1}$, which cannot both be true or false). We migrate this semantic framework into classical FOL with additional constraints, enabling structured multi-perspective reasoning that captures both complex semantics and deep abstract logical relations.
Motivated by agent-based paradigms Hong et al. (2023), we propose LogicAgent, a semiotic-square-guided reasoning framework that automates multi-perspective deduction through a three-stage pipeline: (1) Semantic Structuring Stage constructs a semiotic square to generate perspective variants of a proposition, including its contradiction and contrary, laying the foundation for multi-perspective reasoning and reflection. (2) Logical Reasoning Stage formalizes the contextual premises and performs symbolic deduction along both the original and contradiction paths. (3) Reflective Verification Stage assesses the reasoning trajectory through logic-aware reflection and revises conclusions when inconsistencies arise. This design enables LogicAgent to emulate human-like reasoning while systematically addressing semantic ambiguity and logical complexity.
To address the limitations of existing datasets in evaluating both semantic and logical complexity, we introduce RepublicQA, a benchmark derived from classical philosophical concepts. RepublicQA is manually annotated and cross-validated by multiple annotators to ensure high quality and consistency. Unlike existing reasoning benchmarks such as ProofWriter Tafjord et al. (2020), ProntoQA Saparov & He (2022), FOLIO Han et al. (2022), and ProverQA Qi et al. (2025), which rely on simple templates and semantically straightforward content, RepublicQA is designed to capture semantic complexity through abstract philosophical concepts and logical complexity by reasoning over contradictory or opposing positions. It consistently exceeds existing benchmarks in reading difficulty, lexical diversity, and long-span expressiveness. These properties make RepublicQA a unique testbed for reasoning that requires sensitivity to linguistic subtleties and deep logical relationships.
@@ 2 Preliminaries
Greimas’ Semiotic Square. The Greimas’ Semiotic Square Greimas et al. (1982) is a foundational construct in structuralist semantics that organizes conceptual contraries and contradictions into a four-element structure, enabling fine-grained reasoning over meaning, opposition, and implication. In our work, we migrate this semantic structure into the setting of classical FOL, with additional constraints to ensure logical soundness. Specifically, we introduce an existential import check to avoid vacuous truth from material implication, and extend the evaluation space from binary {True, False} to a three-valued scheme {True, False, Uncertain}, which better reflects reasoning under ambiguity.
@@
Semantic Implication Theorem. If $\color[rgb]{0,0,0.8}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0.8}S_{1}$} and $\color[rgb]{0,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.4,0.4}S_{2}$} are contraries, then within the semiotic square the following semantic implications hold:
@@
Figure 3: Overview of the LogicAgent framework. The agent processes a natural language proposition through three modular reasoning stages. (1) Semantic Structuring Stage constructs a Greimas’ Semiotic Square, generating four interrelated propositions: the primary proposition $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}S_{1}$}, its contradiction $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\lnot S_{1}$}, the contrary $\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}S_{2}$}, and the contradiction of the contrary $\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\lnot S_{2}$}. These are verified for FOL-consistency using a CFG-based parser. (2) Logical Reasoning Stage transforms the premises into FOL, plans deductive steps for each proposition, and performs symbolic reasoning to evaluate their answers. (3) Reflective Verification Stage adjudicates the final judgment via three procedures: Direct Resolution, applied when $\color[rgb]{0,0,0.8}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0.8}S_{1}$} and $\color[rgb]{0.7,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.7,0,0}\lnot S_{1}$} offer a contradictory answer; Quick Reflection, used when either $\color[rgb]{0,0,0.8}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0.8}S_{1}$} or $\color[rgb]{0.7,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.7,0,0}\lnot S_{1}$} is uncertain; and Deep Reflection, used when both $\color[rgb]{0,0,0.8}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0.8}S_{1}$} and $\color[rgb]{0.7,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.7,0,0}\lnot S_{1}$} yield the same value, requiring further validation through the semiotic implication relations involving $\color[rgb]{0,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.4,0.4}S_{2}$} and $\color[rgb]{0.8,0.4,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0.4,0}\lnot S_{2}$}.
@@ 3.2 Semantic Structuring Stage
Contradictory Construction.
Given a natural-language proposition $S_{1}$, we first formalize it into a FOL expression. We then negate the entire formula to obtain $\lnot S_{1}$ and simplify it using standard equivalences (quantifier negation, De Morgan, implication, bi-implication), as summarized in Table 1 (column $S_{1}$ (simplified)”). Finally, the simplified form is mapped back into natural language, ensuring that $\lnot S_{1}$ is both syntactically valid in FOL and semantically a strict negation of $S_{1}$.
Contrary Construction.
We adopt the classical definition of contrariety: $S_{1}$ and $S_{2}$ cannot both be true but may both be false. To block vacuous truth, we enforce non-emptiness and relevance constraints; when unmet, the evaluation defaults to Uncertain. The construction follows three paths:
  • Strict contraries: For canonical templates that always yield valid contraries (Table 1, “Strict”), we apply symbolic rules directly.
  • Conditional contraries: For templates requiring constraints to avoid vacuity (Table 1, “Conditional”), we first generate candidate contraries $S_{2}$ (via rules/LLM). Each undergoes an existential import check against premises $P$ to ensure non-empty domains, satisfiable antecedents, and instantiable variables. Valid candidates are then tested for contrariety with $S_{1}$ and contextual relevance (e.g., accept $A\!\to\!\lnot B$ only if $\mathrm{Sat}_{P}(A)$ holds).
  • Beyond coverage: The six rules in Table 1 cover common logical forms. For other cases (nested or domain-specific), we rely on model-assisted generation with self-supervised validation.
Validation and Verification.
All candidate propositions ($S_{1}$, $\lnot S_{1}$, $S_{2}$, $\lnot S_{2}$) are validated through a three-stage pipeline, and only those passing all stages are retained for downstream reasoning:
@@
# $S_{1}$ $\lnot S_{1}$ $S_{2}$ Type Constraint
1 $\forall x\,\varphi(x)$ $\lnot\forall x\,\varphi(x)\ \equiv\ \exists x\,\lnot\varphi(x)$ $\forall x\,\lnot\varphi(x)$ Strict N/A
2 $A\wedge B$ $\lnot(A\wedge B)\ \equiv\ \lnot A\vee\lnot B$ $A\wedge\lnot B$ Strict N/A
3 $A\leftrightarrow B$ $\lnot(A\leftrightarrow B)\ \equiv\ A\oplus B$ $A\leftrightarrow\lnot B$ Strict N/A
4 $\exists x\,\varphi(x)$ $\lnot\exists x\,\varphi(x)\ \equiv\ \forall x\,\lnot\varphi(x)$ $\exists x\,\lnot\varphi(x)$ Conditional $D\neq\emptyset$ (non-empty domain)
5 $A\rightarrow B$ $\lnot(A\rightarrow B)\ \equiv\ A\wedge\lnot B$ $A\rightarrow\lnot B$ Conditional $\text{Sat}(A)$ (antecedent satisfiable)
6 $A\vee B$ $\lnot(A\vee B)\ \equiv\ \lnot A\wedge\lnot B$ $A\vee\lnot B$ Conditional $A=\mathbf{F}$ (antecedent exclusion)

### v5

@@ Abstract
Logical reasoning is a fundamental capability of large language models (LLMs). However, existing studies largely overlook the interplay between logical complexity and semantic complexity, resulting in methods that struggle to address challenging scenarios involving abstract propositions, ambiguous contexts, and conflicting stances, which are central to human reasoning. We propose LogicAgent, a semiotic-square–guided framework that jointly addresses these two axes of difficulty. The semiotic square provides a principled structure for multi-perspective semantic analysis, and LogicAgent integrates automated deduction with reflective verification to manage logical complexity across deeper reasoning chains. To support evaluation under these conditions, we introduce RepublicQA, a benchmark that couples semantic complexity with logical depth. RepublicQA reaches college-level semantic difficulty (FKGL 11.94), contains philosophically grounded abstract propositions with systematically constructed contrary and contradictory forms, and offers the most semantically rich setting for assessing logical reasoning in LLMs. Experiments demonstrate that LogicAgent achieves state-of-the-art performance on RepublicQA, with a 6.25% average gain over strong baselines, and generalizes effectively to mainstream logical reasoning benchmarks including ProntoQA, ProofWriter, FOLIO, and ProverQA, achieving an additional 7.05% average gain. These results highlight the strong effectiveness of our semiotic-grounded multi-perspective reasoning in boosting LLMs’ logical performance. Code is available at https://github.com/AI4SS/Logic-Agent.
@@ 1 Introduction
Logical reasoning (Smith, 2003) plays a central role in human cognition, enabling structured transitions from ambiguous inputs to definitive conclusions. In AI (Cohen et al., 2020; Vaswani et al., 2017), it underpins tasks such as commonsense reasoning (Wang et al., 2024), mathematical proof (Wang et al., 2023; Eisner et al., 2024; Gao et al., 2023), and philosophical thinking (Paul, 2013). However, robust logical reasoning in natural language remains challenging due to (1) semantic complexity (Tuggy, 1993), where expressions admit multiple interpretations or surface forms, and (2) logical complexity (Gibson, 1998), which requires reasoning over contextual premises and semantic interactions (Zhang et al., 2020; Ding et al., 2024).
Recent approaches can be grouped into three categories: (1) Linear Reasoning (LR) methods, such as Naive Prompting and Chain-of-Thought (Wei et al., 2022); (2) Aggregative Reasoning (AR) approaches (Ryu et al., 2024; Zhang et al., 2023; Yao et al., 2023; Sun et al., 2024) that combine multiple reasoning trajectories; and (3) Symbolic Reasoning (SR) frameworks (Yang et al., 2023; Xu et al., 2024a; b) that integrate LLMs (Achiam et al., 2023; Guo et al., 2025; Zhao et al., 2023b) with explicit symbolic modules (Pan et al., 2023). Although effective, these methods remain centered on logical structure and give limited attention to semantic complexity, often assuming clean predicates and unambiguous contexts. This neglects how abstraction, conflicting stances, and contextual ambiguity interact with logical reasoning, limiting performance when semantic and logical complexity jointly shape reasoning (Lago, 2009).
@@
To capture the complex semantics and deep logical relations, we draw inspiration from Greimas’ Semiotic Square (Greimas et al., 1982; Greimas, 1987; 1988), a structuralist framework that extends binary oppositions into a four-part structure. It encompasses both contraries (e.g., $S_{1}$ vs. $S_{2}$, which cannot both be true but may both be false under non-empty domains) and contradictions (e.g., $S_{1}$ vs. $\lnot S_{1}$, which cannot both be true or false). We migrate this semantic framework into classical FOL with additional constraints, enabling structured multi-perspective reasoning that captures both complex semantics and deep abstract logical relations.
Motivated by agent-based paradigms (Hong et al., 2023), we propose LogicAgent, a semiotic-square-guided reasoning framework that automates multi-perspective deduction through a three-stage pipeline: (1) Semantic Structuring Stage constructs a semiotic square to generate perspective variants of a proposition, including its contradiction and contrary, laying the foundation for multi-perspective reasoning and reflection. (2) Logical Reasoning Stage formalizes the contextual premises and performs symbolic deduction along both the original and contradiction paths. (3) Reflective Verification Stage assesses the reasoning trajectory through logic-aware reflection and revises conclusions when inconsistencies arise. This design enables LogicAgent to emulate human-like reasoning while systematically addressing semantic ambiguity and logical complexity.
To rigorously evaluate how semantic complexity interacts with logical reasoning, we introduce RepublicQA, a benchmark grounded in classical philosophical concepts and annotated through multi-stage, cross-validated human review. Existing reasoning benchmarks such as ProofWriter (Tafjord et al., 2020), ProntoQA (Saparov and He, 2022), FOLIO (Han et al., 2022), and ProverQA (Qi et al., 2025) are largely template-based and focus primarily on logical structure, offering limited semantic depth and little coverage of the ways semantic ambiguity, abstraction, or opposing stances influence logical inference. In contrast, RepublicQA captures semantic complexity through abstract philosophical content and logical complexity through systematically organized contrary and contradictory relations. It also exceeds existing benchmarks across all five semantic complexity indicators, reaching a college-level reading difficulty (FKGL = 11.94) while maintaining the same level of logical reasoning rigor required by prior datasets.
@@ 2 Preliminaries
Greimas’ Semiotic Square. The Greimas’ Semiotic Square (Greimas et al., 1982) is a foundational construct in structuralist semantics that organizes conceptual contraries and contradictions into a four-element structure, enabling fine-grained reasoning over meaning, opposition, and implication. In our work, we migrate this semantic structure into the setting of classical FOL, with additional constraints to ensure logical soundness. Specifically, we introduce an existential import check to avoid vacuous truth from material implication, and extend the evaluation space from binary {True, False} to a three-valued scheme {True, False, Uncertain}, which better reflects reasoning under ambiguity.
@@
Theorem 1 (Semantic Implication Theorem).
If $S_{1}$ and $S_{2}$ are contraries, then within the semiotic square the following semantic implications hold:
@@ 3.1 Task Definition
Figure 3: Overview of the LogicAgent framework. The agent processes a natural language proposition through three stages. (1) Semantic Structuring Stage constructs a Greimas’ Semiotic Square, generating four interrelated propositions: the primary proposition $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}S_{1}$}, its contradiction $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\lnot S_{1}$}, the contrary $\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}S_{2}$}, and the contradiction of the contrary $\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\lnot S_{2}$}. These are verified for FOL-consistency using a CFG-based parser. (2) Logical Reasoning Stage transforms the premises into FOL, plans deductive steps for each proposition, and performs symbolic reasoning to evaluate their answers. (3) Reflective Verification Stage adjudicates the final judgment via three procedures: Direct Resolution, applied when $\color[rgb]{0,0,0.8}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0.8}S_{1}$} and $\color[rgb]{0.7,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.7,0,0}\lnot S_{1}$} offer a contradictory answer; Quick Reflection, used when either $\color[rgb]{0,0,0.8}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0.8}S_{1}$} or $\color[rgb]{0.7,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.7,0,0}\lnot S_{1}$} is uncertain; and Deep Reflection, used when both $\color[rgb]{0,0,0.8}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0.8}S_{1}$} and $\color[rgb]{0.7,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.7,0,0}\lnot S_{1}$} yield the same value, requiring further validation through the semiotic implication relations involving $\color[rgb]{0,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.4,0.4}S_{2}$} and $\color[rgb]{0.8,0.4,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0.4,0}\lnot S_{2}$}.
@@ 3.2 Semantic Structuring Stage
Contradictory Construction. Given a natural-language proposition $S_{1}$, we first formalize it into a FOL expression. We then negate the entire formula to obtain $\lnot S_{1}$ and simplify it using standard equivalences (quantifier negation, De Morgan, implication, bi-implication), as summarized in Table 1 (column $S_{1}$ (simplified)”). Finally, the simplified form is mapped back into natural language, ensuring that $\lnot S_{1}$ is both syntactically valid in FOL and semantically a strict negation of $S_{1}$.
Definition 1 (Existential Import Check).
Let $P$ be the set of premises defining a model $\mathcal{M}_{P}=(D_{P},I_{P})$, where $D_{P}$ is the domain of discourse and $I_{P}$ the interpretation function. For a candidate formula $\phi$, its existential import under $P$ holds, written $\mathrm{EIC}_{P}(\phi)=\mathbf{T}$, iff
$\exists\eta:\text{Free}(\phi)\!\to\!D_{P}\ \text{such that}\ \mathcal{M}_{P},\eta\models\text{Ante}(\phi),$
where $\text{Ante}(\phi)$ denotes the antecedent or quantifier scope of $\phi$. Otherwise $\mathrm{EIC}_{P}(\phi)=\mathbf{F}$, indicating that $\phi$ is vacuous (e.g., empty domain or unsatisfiable antecedent).
Lemma 1 (Soundness of Conditional Contrariety).
A candidate pair $({\color[rgb]{0,0,0.8}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0.8}S_{1}},{\color[rgb]{0,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.4,0.4}S_{2}})$ generated under a rule $r$ in Table 1 is a valid contrary pair under $P$ iff both satisfy $\mathrm{EIC}_{P}({\color[rgb]{0,0,0.8}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0.8}S_{1}})=\mathrm{EIC}_{P}({\color[rgb]{0,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.4,0.4}S_{2}})=\mathbf{T}$ and $\mathcal{M}_{P}\models\lnot({\color[rgb]{0,0,0.8}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0.8}S_{1}}\wedge{\color[rgb]{0,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.4,0.4}S_{2}})$. Pairs failing either condition are excluded from the contrary set.
This establishes that only non-vacuous and mutually unsatisfiable pairs are retained, ensuring that the migration from semiotic structure to FOL preserves logical soundness.
Contrary Construction. We adopt the classical definition of contrariety: $S_{1}$ and $S_{2}$ cannot both be true but may both be false. Table 1 summarizes six unified rules for constructing contraries and contradictories. For strict forms, symbolic transformation directly yields valid contraries. For conditional forms, candidates $S_{2}$ are first generated (via rules or LLM transformation) and then verified by the existential import check (Definition 1) to ensure non-vacuous quantifiers and satisfiable antecedents. Only pairs satisfying $\mathrm{EIC}_{P}({\color[rgb]{0,0,0.8}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0.8}S_{1}})=\mathrm{EIC}_{P}({\color[rgb]{0,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.4,0.4}S_{2}})=\mathbf{T}$ and $\mathcal{M}_{P}\models\lnot({\color[rgb]{0,0,0.8}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0.8}S_{1}}\wedge{\color[rgb]{0,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.4,0.4}S_{2}})$ are retained as valid contraries (Lemma 1). For structures beyond these six templates, model-assisted generation with self-supervised validation re-applies Definition 1 to filter logically sound pairs.
Validation and Verification. All candidate propositions ($S_{1}$, $\lnot S_{1}$, $S_{2}$, $\lnot S_{2}$) are validated through a three-stage pipeline, and only those passing all stages are retained for downstream reasoning:
@@
# $S_{1}$ $\lnot S_{1}$ $S_{2}$ Type Constraint EIC Condition
1 $\forall x\,\varphi(x)$ $\lnot\forall x\,\varphi(x)\ \equiv\ \exists x\,\lnot\varphi(x)$ $\forall x\,\lnot\varphi(x)$ Strict N/A Always holds
2 $A\wedge B$ $\lnot(A\wedge B)\ \equiv\ \lnot A\vee\lnot B$ $A\wedge\lnot B$ Strict N/A Always holds
3 $A\leftrightarrow B$ $\lnot(A\leftrightarrow B)\ \equiv\ A\oplus B$ $A\leftrightarrow\lnot B$ Strict N/A Always holds
4 $\exists x\,\varphi(x)$ $\lnot\exists x\,\varphi(x)\ \equiv\ \forall x\,\lnot\varphi(x)$ $\exists x\,\lnot\varphi(x)$ Conditional $D=\emptyset$ $\mathrm{EIC}_{P}(\varphi(x))=\mathbf{F}$
5 $A\rightarrow B$ $\lnot(A\rightarrow B)\ \equiv\ A\wedge\lnot B$ $A\rightarrow\lnot B$ Conditional $\text{Sat}(A)$ $\mathrm{EIC}_{P}(A)=\mathbf{T}$
6 $A\vee B$ $\lnot(A\vee B)\ \equiv\ \lnot A\wedge\lnot B$ $A\vee\lnot B$ Conditional $A=\mathbf{F}$ $\mathrm{EIC}_{P}(B)=\mathbf{T}$